What is the length of a side of a square if the area and perimeter are numerically equal?: What is the length of a side of a square if the area and perimeter are numerically equal?
Answer by checkley77(3654)

4*x=x^2
x^2-4x=0
x(x-4)=0
x=4 answer.
Proof:
4*4=4^2
16=16